right wrong
Pick an answer to see how members did
Log this miss in the free GMAT error log
Get the error logGet one like this every day: the free 60-day plan
Start the plan(A) 1/495
(B) 1/165
(C) 1/81
(D) 1/3
(E) 1/2
Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.
See classesLearn live with an expert or move at your own pace. Every option includes the complete TTP study system.
Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.
with Josh Braslow, 20+ years in test prep
Self-paced EA prep. Study on your schedule.
with
Complete access from day one. Study on your schedule.
right wrong
Pick an answer to see how members did
Log this miss in the free GMAT error log
Get the error logGet one like this every day: the free 60-day plan
Start the planSo the situation is that there are 4 marbles each of red, green, and blue colors.BTGmoderatorDC wrote:A jar contains 12 marbles consisting of an equal number of red, green, and blue marbles. Four marbles are removed from the jar and discarded. What is the probability that only two colors will remain in the jar after the four marbles have been removed?
(A) 1/495
(B) 1/165
(C) 1/81
(D) 1/3
(E) 1/2
55.9 should be 495Jay@ManhattanReview wrote:So the situation is that there are 4 marbles each of red, green, and blue colors.BTGmoderatorDC wrote:A jar contains 12 marbles consisting of an equal number of red, green, and blue marbles. Four marbles are removed from the jar and discarded. What is the probability that only two colors will remain in the jar after the four marbles have been removed?
(A) 1/495
(B) 1/165
(C) 1/81
(D) 1/3
(E) 1/2
We have to find out the probability that only two colors will remain in the jar after the four marbles have been removed and discarded.
Say the 4 marbles removed are red, thus,
the probability that only green, and blue colors will remain in the jar after 4 red marbles have been removed = 4C4 / 12C4 = 1/12C4 = (1.2.3.4)/(12.11.10.9) = 1/(55.9)
In the same way, the 4 marbles removed can be green or blue, thus,
the probability that only two colors will remain in the jar after the four marbles have been removed and discarded
= 3*[1/(55.9)] = 1/165
The correct answer: B
Hope this helps!
-Jay
_________________
Manhattan Review Test Prep
Locations: Manhattan Review Chennai | Free GMAT Practice Test | GRE Prep Hyderabad | Jayanagar GRE Coaching | and many more...
Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
For only 2 colors to remain after 4 marbles have been removed, the 4 selected marbles must be of the same color.BTGmoderatorDC wrote:A jar contains 12 marbles consisting of an equal number of red, green, and blue marbles. Four marbles are removed from the jar and discarded. What is the probability that only two colors will remain in the jar after the four marbles have been removed?
(A) 1/495
(B) 1/165
(C) 1/81
(D) 1/3
(E) 1/2
My day for nagging corrections I guess. 1/11 should be 3/11GMATGuruNY wrote:For only 2 colors to remain after 4 marbles have been removed, the 4 selected marbles must be of the same color.BTGmoderatorDC wrote:A jar contains 12 marbles consisting of an equal number of red, green, and blue marbles. Four marbles are removed from the jar and discarded. What is the probability that only two colors will remain in the jar after the four marbles have been removed?
(A) 1/495
(B) 1/165
(C) 1/81
(D) 1/3
(E) 1/2
The first selected marble can be of ANY COLOR.
A good outcome will be yielded if the second, third and fourth marbles are of the SAME COLOR as the first.
P(2nd marble is of the same color as the first) = 3/11. (Of the 11 marbles that remain after the removal of the 1st marble, 3 are of the same color as the first.)
P(3rd marble is of the same color as the first 2) = 2/10. (Of the 10 remaining marbles, 2 are of the same color as the first two.)
P(4th marble is of the same color as the first 3) = 1/9. (Of the 9 remaining marbles, 1 is of the same color as the first three.)
To combine these probabilities, we multiply:
1/11 * 2/10 * 1/9 = 1/165.
The correct answer is B.
Fixed!regor60 wrote:My day for nagging corrections I guess. 1/11 should be 3/11
There are 4 red, 4 green, and 4 blue marbles in the jar.BTGmoderatorDC wrote:A jar contains 12 marbles consisting of an equal number of red, green, and blue marbles. Four marbles are removed from the jar and discarded. What is the probability that only two colors will remain in the jar after the four marbles have been removed?
(A) 1/495
(B) 1/165
(C) 1/81
(D) 1/3
(E) 1/2
